Given dy/dx=y²/25 and y = 5 when x = 0 Find x when y =2


1
Expert's answer
2021-05-24T18:40:24-0400

Let us solve the differential equation "\frac{dy}{dx}=\frac{y^2}{25}":


"\frac{dy}{y^2}=\frac{dx}{25}"


"\int\frac{dy}{y^2}=\int\frac{dx}{25}"


"-\frac{1}{y}=\frac{x}{25}+C"


Since "y = 5" when "x = 0", we have that "C=-\frac{1}{5}."


It follows that "-\frac{1}{y}=\frac{x}{25}-\frac{1}{5}."


Let us find "x" when "y =2:"


"-\frac{1}{2}=\frac{x}{25}-\frac{1}{5}"


"\frac{x}{25}=\frac{1}{5}-\frac{1}{2}=-\frac{3}{10}"


We conclude that "x=-\frac{15}{2}."